Laser Subcommittee Meeting 2019-04-08

  • Attendees: FletcherB, NickB, BlakeS
  • LC1290 Big Red
    • Swapped to the new (red) head. The tip of the nozzle was drilled out to 5mm to accommodate the beam of the laser. The z-probe mount was expanded to 25mm to fit on the new head. Right now it's a little difficult to disassemble the head for lens cleaning so best to leave it unless you've got some extra time on your hands.
    • The new z offset is 15.5mm from the probe and 23.5mm from the tip of the head. The laser is cutting 45-55mm/s on 3mm MDF comfortably.
    • Bed levelled
  • Kellogg
    • Following on from our last session with Kellogg Blake found a diagnostic menu that indicates that one of the water pump interlocks is showing as active, we'll need to either disable it temporarily or figure out how to connect it to the chiller before testing can continue. There's no jack on the back of the cutter for the interlock (unlike Big Red).
  • subcommittee/laser-minutes-20190408.txt
  • Last modified: 2020/03/12 12:49
  • (external edit)